OpenStudy (anonymous):

What percent of 94 is 23? (1 point)

OpenStudy (anonymous):

These are the choices

OpenStudy (anonymous):



OpenStudy (anonymous):

I dont understand....

OpenStudy (anonymous):

you have to figure out the equation x*94 = 23

OpenStudy (anonymous):

It will give you a decimal answer for 94, which then you can convert to a %

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so x = 23/94 = .244 which is about 24%

OpenStudy (anonymous):

Then to check your answer, do .24 * 93

OpenStudy (anonymous):

hmm sorry, that isn't right. Hold on a second

OpenStudy (anonymous):

you just have to multiply both sides of n/94 = 23/100 by 94. n = 23*94/100 = 21.62